OpenStudy (lovelyanna):

The sum of three consecutive odd integers is –15. What are the numbers? A. {–3, –1, 1} B. {–3, –5, –7} C. {–7, –9, –11} D. {–4, –5, –6} B?

OpenStudy (lovelyanna):

@jigglypuff314

OpenStudy (shamil98):

Three consecutive odd integers, it's pretty self-explanatory.. what odd numbers added together give you -15? look at your answer choices..

OpenStudy (lovelyanna):

I said B

OpenStudy (shamil98):

Consecutive means the next number in line. Consecutive normally would be like 1,2,3,4 Consecutive even would be like 2,4,6,8 Consecutive odd would be like 1,3,5,7.

OpenStudy (shamil98):

Yes, b is correct.

OpenStudy (lovelyanna):

Thanks, but you could have just said if B was correct or not. I was just making sure its correct

OpenStudy (lovelyanna):

@jigglypuff314 Hi :) I have a few math questions, I solved them all, but want to make sure, you busy right now?

OpenStudy (lovelyanna):

make sure if they are correct* my bad

jigglypuff314 (jigglypuff314):

sure :)

OpenStudy (lovelyanna):

Thanks :)

OpenStudy (lovelyanna):

Solve for x. (x-7)/8 = 7/2 A. x = 9 2/7 B. x = 63 C. x = 31.5 D. x = 35 D?

jigglypuff314 (jigglypuff314):

correct :)

OpenStudy (lovelyanna):

Awesome! thanks :) next?

jigglypuff314 (jigglypuff314):

sure :)

OpenStudy (shamil98):

if you just need to check algebra problems, just use: www.wolframalpha.com

OpenStudy (lovelyanna):

Write an equation in slope-intercept form for the line passing through (–1, 10) and (5, 8). A. y = 1/3x + 29/3 B. y = -1/3x - 29/3 C. y = -1/3 + 29/3 D. y = 1/3x - 29/3 I'm not sure about this one, is it A?

jigglypuff314 (jigglypuff314):

first find slope... it would be negative I think...

OpenStudy (lovelyanna):

Oh, so it would be -1/3x, and addition?

jigglypuff314 (jigglypuff314):

yup but not always addition, but in this case, yes :)

OpenStudy (lovelyanna):

I would say C then?

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (lovelyanna):

Great, thanks! =)

jigglypuff314 (jigglypuff314):

I gtg :/

OpenStudy (lovelyanna):

Oh, ok. See you later :) when will you be back? or you wont..?

OpenStudy (anonymous):

The sum of three consecutive odd integers is –15. What are the numbers? A. {–3, –1, 1} B. {–3, –5, –7} C. {–7, –9, –11} D. {–4, –5, –6} Let x be the middle number. Solve the following for x: (x-2)+x+{x+2}=-15 x=-5 The numbers from low to high are: {-7,-5,-3}

OpenStudy (lovelyanna):

Thanks, but this isn't the question I need help with

OpenStudy (anonymous):

Is is the following: Write an equation in slope-intercept form for the line passing through (–1, 10) and (5, 8).

OpenStudy (lovelyanna):

No, I tagged you on the question :)